linear expansion is defined from the equation

6. Linear expansion (a) is defined form the equation (a) L-Lo/a LoAt (b) L=Lo-a LoAt (c) L-Lot a LoAt (d) None
Answer #1

linear expansion is defined from the equation given in OPTION C.

\frac{\Delta L}{L_{0}} = \alpha \Delta T

where \Delta L = L - L_0

now L - L_0 = L_0 \alpha \Delta T

L = L_0 + L_0 \alpha \Delta T <<<<<<<<Option C